OpenStudy (anonymous):

a car slows down from a speed of 25m/s to rest in 5.00s. How far did travel in that time?

OpenStudy (anonymous):

initial speed=u=25m/s,final speed=v=0m/s,time=T=5sec. v=u+at 0=25+a*5 -25=a*5 a=-5metre per second square using third law:- 0-625=2*(-5)*S -625= -10*S S= -625/-10 S=62.5m.

OpenStudy (safari321):

STUDY THE MOTION IN A STRAIGHT LINE SECTION OF UR TEXT BOOK IF YOU HAVE ON :l

OpenStudy (safari321):

:\